OpenStudy (anonymous):

Need help for test tomorrow solve please Solve the system by substitution or elimination. Write the answer from the smallest x value to the largest. ie (-2, 5) (3, -4) (7, 2) -x - y + 3 = 0 x2 - 4x + 7 = y

OpenStudy (steve816):

I'll give you help! Why don't you go to sleep so you can focus during the test :)

OpenStudy (anonymous):

Thank you

OpenStudy (steve816):

No problem!

OpenStudy (anonymous):

Can you also answer this Solve the system by substitution or elimination. Write the answer from the smallest x value to the largest. ie (-2, 5) (3, -4) (7, 2) x - 2y = 0 3x - y2 = 0

OpenStudy (lynfran):

multiply eq1 by -3 we get -3x+6y=0 3x-y^2=0 add the 2 equations 6y-y^2=0 y(6-y)=0 y=0 and y=6 sub y=0 and or y=6 in eq1 x-2y=0 x-2(6)=0 x=12 when x=12 y=6 (12, 6) x-2(0)=0 x=0 when x=0 y=0 (0,0)

OpenStudy (anonymous):

Thank you can you also solve the first equation

OpenStudy (lynfran):

ok

OpenStudy (steve816):

Wow, great explanation @LynFran You are really great with math :)

OpenStudy (lynfran):

-x-y+3=0 x^2-4x+7=y rewrite eq1 to -x+3=y now equate the 2 equations we get x^2-4x+7=-x+3 x^2-4x+x=3-7 x^2-3x=-4 (x- 1.5)^2=-4+2.25 (x-1.5)^2=-1.75 ....................please check ur equation to make sure its correctly type..im getting nowhere with this..

OpenStudy (anonymous):

-x - y + 3 = 0 x2 - 4x + 7 = y This is it

OpenStudy (anonymous):

Can u help me on this You are upset with your satellite TV provider and choose to build your own parabolic satellite dish. You want the dish to be 12 inches across. You want the maximum height of the dish to be 7 feet. What is the value of "a" in the parabolic equation for the dish?